# Computing with D-Algebraic Sequences

### v1

@@ 4.2 $C^{2}$-Finite sequences
It suffices to construct a dynamical system in the form of (7) where $\mu=(1,\ldots,1)^{T}$ and $\mathbf{E}_{\mathbf{A}}(\sigma(\mathbf{w}))=\boldsymbol{0}$. This is straightforward from the formal definition of a $C^{2}$-finite sequence. Let $(s(n))_{n}$ be $C^{2}$-finite of order $l$ with $C$-finite coefficients $c_{l}(n),\ldots,c_{0}(n)$, of order $r_{l},\ldots,r_{0}$, respectively. For simplicity, we assume $N=0$. From the defining equation of the coefficient $c_{j}(n)$, we can write
$c_{j}(n+r_{i})=\alpha_{0,j}\,c_{i}(n)+\cdots+\alpha_{r_{j}-1,j}\,c_{j}(n+l-1),$ (28)
where $\alpha_{i,j}\in\mathbb{K}$, $i=0,\ldots,r_{j}-1,j=0,\ldots,l$. Similarly, from the equation of $s(n)$, we have
$s(n+l)=-\frac{1}{c_{l}(n)}\left(c_{0}(n)\,s(n)+\cdots+c_{l-1}(n)\,s(n+l-1)\right).$ (29)
Thus, following the construction from Section 3.1, we deduce that each $c_{j}(n)$ adds $r_{j}$ equations to the system, which leads to a rational dynamical system of dimension $l+\sum_{j=0}^{l}r_{j}$. ∎
In Theorem 4.5, $N$ is known to exist for non-degenerate $C$-finite sequences [BM76]. In the general case, however, the proper definition of a $C^{2}$-finite sequence is subject to answering the Skolem problem [OW12] for the leading $C$-finite term coefficient. Note that this constraint highly depends on the choice of initial values: the sequence $((-1)^{n}-1)_{n}$ satisfies the recurrence $s(n+2)-s(n)=0$ and has infinitely many zeros; however, the sequence $(2\,(-1)^{n}-1)_{n}$ satisfies the same recurrence equation but has no zero. In the following example, we use our algorithm to compute an algebraic difference equation satisfied by a $C^{2}$-finite sequence for which two not-necessarily-equal solutions of $s(n+2)=s(n)$ appear within the coefficients of its equation.
@@
$u(n)\,s(n)+2\,s(n+1)+v(n)\,s(n+2)=0,$ (30)
where $(u(n))_{n}$ and $(v(n))_{n}$ are two not-necessarily identical solutions of $s(n+2)-s(n)=0$. This includes the result of [JPNP23, Example 4.1]. The corresponding dynamical system writes:
$\begin{cases}\sigma(w_{1})=w_{2}\\ \sigma(w_{2})=-\frac{w_{5}\,w_{1}+2\,w_{2}}{w_{3}}\\ \sigma(w_{3})=w_{4}\\ \sigma(w_{4})=w_{3}\\ \sigma(w_{5})=w_{6}\\ \sigma(w_{6})=w_{5}\\ z=w_{1}\end{cases}.$ (31)
The resulting $6$th-order algebraic difference equation is l.h.s. and may be written recursively as follows:
$s(n+6)=\frac{s(n+5)s(n+4)s(n)+s(n+4)s(n+3)s(n+2)-s(n+4)^{2}s(n+1)-s(n+5)s(n+2)^{2}}{s(n+3)s(n)-s(n+2)s(n+1)}.$ (32)
$\blacksquare$
Although we obtained an l.h.s. algebraic difference equation in (32), the underlying adaptation of Algorithm 1 does not guarantee such an output. A bound for the order of the desired l.h.s. equation is provided in [CDBMP23] but without an algorithm.
@@ 5 Subsequences
Recall that a subsequence of a sequence $(s(n))_{n}$ is a sequence $(t(n))_{n}$ such that $t(n)=s(\varphi(n))$, where $\varphi$ is a strictly increasing embedding of $\mathbb{N}$. Thus, to find a difference polynomial $p$ that vanishes at $(t(n))_{n}$, we need an endomorphism $\tilde{\sigma}$ that sends $s(\varphi(n))$ to $s(\varphi(n+1))$, i.e., the minimal gap between the orders of two difference monomials in $p$ is $\delta_{\tilde{\sigma}}=\min_{n\in\mathbb{N}}\{\varphi(n+1)-\varphi(n)\}$ and not the usual $n+1-n=1$. This section focuses on the case where $\varphi(n)=dn$. This is the $n$th term formula of an arithmetic progression of common difference $d$ with initial term 00. What makes this case rather natural is that all order gaps are multiple of the common difference $d$, and here we have $\delta_{\tilde{\sigma}}=d$. This implies that the normal form of $p$ can be seen as follows:
@@
The desired endomorphism is thus defined by $\tilde{\sigma}=\sigma^{d}$.
@@
Our aim is to eliminate the difference indeterminates $y_{j}$, $j>1$. To do so, we build a radical-rational dynamical system and apply Algorithm 1. This will yield a univariate difference polynomial in $y_{1}$ that has $(s(dn))_{n\in\mathbb{N}}$ as a zero. Let $m$ be the degree of $\sigma^{r}(x)$ in $p$ and $R_{p}$ be the rational expression obtained by solving $p=0$ for $(\sigma^{r}(x))^{m}$. Let $y_{i,j},i=1,\ldots,d$, $j=0,\ldots,r_{1}$ be new indeterminates such that
@@
The proof reduces to determining the dimension of the dynamical system obtained in the proof of Theorem 5.1. Indeed, we have
@@
The recurrence equation of the subsequence need not be written in the same difference ring as the equation of the original sequence. In Theorem 5.1, both $(s(n))_{n\in\mathbb{N}}$ and $(s(dn))_{n\in\mathbb{N}}$ are zeros of the constructed difference polynomial of order $dr$ in $\mathcal{D}_{\sigma}(\mathbb{K},x)$, but $(s(n))_{n}$ is generally not a zero of the corresponding difference polynomial of order $r$ in $\mathcal{D}_{\sigma^{d}}(\mathbb{K},x)$. This is illustrated in the next example.
@@
In Section 4.1, Example 4.4, we used the implementation from [TTW24] to convert the holonomic equation of Catalan numbers $(C(n))_{n\in\mathbb{N}}$ to the recursion
@@
One verifies that $(C(n))_{n}$ is not a solution of (51). Note that higher-order equations may be obtained here. Indeed, $(C(3n))_{n}$ is holonomic of order $1$ and degree $3$ and therefore satisfies an l.h.s. algebraic difference equation of order at most $4$ by Theorem 4.2. Our Gröbner bases method from [TTW24] yields an equation of order $3$blacksquare
Returning to our introductory motivation—the acceleration of sequences in numerical analysis—we have demonstrated a method for deriving equations for subsequences of D-algebraic sequences, indexed by arithmetic progressions. The method may be generalized to a broader class of subsequences. This work suggests a potential bridge between difference algebra and classical acceleration techniques (see [BDGB83]). Careful selection of subsequences based on these computations may lead to the development of new effective methods for accelerating convergence.
@@ Acknowledgments
The author gratefully acknowledges Filip Mazowiecki and James Worrell for stimulating discussions regarding [CDBMP23]. This work was supported by UKRI Frontier Research Grant EP/X033813/1.

### v2

@@ 4.2 $C^{2}$-Finite sequences
This is a consequence of the correctness of Algorithm˜2, which might be detailed as follows.
  • Each $c_{j}(n+k),\,0\leq j\leq l,$ and $0\leq k<r_{j}$, is eliminated in step 2.1.1. of Algorithm˜2.
  • This elimination enables the replacement of all appearances of $c_{j}(n+k)$ from the $(q_{j})$, $(p)$, and all $c_{i}(n+k_{i})$ that are already eliminated.
  • Therefore, when the loop in step $2$ ends, each $c_{j}(n+k)$’s on the right hand side of each $C$-finite equation $(q_{j})$ is a rational expression of shifts of $s(n)$, unless all $r_{j}\geq 1$, in which case $c_{l}(n+r_{l}-1)$ does not need to be expressed as a rational expression of shifts of $s(n)$.
  • The number of shifts applied to $(p)$ is exactly $r=l+\sum_{j=0}^{l}r_{j}$, with $r_{l}$ potentially updated in step $1$. In this shift of $(p)$, one of the factors is a rationalizing difference polynomial in which $s(n+r)$ appears linearly.
Remark 4.11.
  • As explained in the first-order case, in Algorithm˜2, when some $C$-finite term coefficients are nonzero constants, i.e, $r_{j}=0$ and $c_{j}(n)\neq 0$, all $c_{j}(n+k),\,0\leq k<r_{j}$ have to be expressed as rational expressions of shifts of $s(n)$, and in this case the order of the output is exactly $l+\sum_{j=0}^{l}r_{j}$, where $r_{l}$ is not updated in step 1.
  • The arrangement of the $C$-finite term coefficients in the input is irrelevant for the algorithm: Interchanging the indices $i$ and $j>i$ to eliminate $c_{j}$ before $c_{i}$ in the list of $C$-finite coefficients does not affect the correctness of the algorithm.
  • Theorem˜4.10 is also another proof of Theorem˜4.5 since every holonomic sequence is $C^{2}$-finite as every polynomial satisfies a $C$-finite equation.
  • The elimination procedure in Algorithm 2 is inherently similar to the linear algebra computations detailed in [JPPS20, Section 4]. Specifically, an analysis of the resultant approach (Theorem 28 in that paper) reveals that this method invariably yields a rationalizing difference polynomial when applied to $C^{2}$-finite equations. This finding was later explained by Jimenez-Pastor (private communication), after we had presented Theorem˜4.10 to him.
We implemented Algorithm˜2 as CCfiniteToSimpleRatrec. The latter is much more efficient than CCfiniteToDalg. The crucial advantage of Algorithm˜2 is that it does not use Gröbner bases. We also observed that when CCfiniteToDalg completes its computations in a reasonable time, its output corresponds with that of CCfiniteToSimpleRatrec.
The approach used in CCfiniteToDalg requires that $c_{l}(n)\neq 0$ for sufficiently large $n$. This condition is satisfied when $(c_{l}(n))_{n}$ is non-degenerate [BM76]. This means that subsequences of $(c_{l}(n))_{n}$ indexed by arithmetic progressions are not $C$-finite of orders lower than $r_{l}$ [BM76, Met00]. In the general case, however, the proper definition of a $C^{2}$-finite sequence is subject to answering the Skolem problem [OW12] for the leading $C$-finite term coefficient. Note that this constraint highly depends on the choice of initial values: the sequence $((-1)^{n}-1)_{n}$ satisfies the recurrence s(n+2)s(n)=0s(n+2)-s(n)=0 and has infinitely many zeros; however, the sequence $(2\,(-1)^{n}-1)_{n}$ satisfies the same recurrence equation but has no zero. In the following example, we use CCfiniteToSimpleRatrec to compute a rational recursion satisfied by a $C^{2}$-finite sequence for which two not-necessarily-equal solutions of $s(n+2)=s(n)$ appear within the coefficients of its equation.
@@
u(n)s(n)+2s(n+1)+v(n)s(n+2)=0,u(n)\,s(n)+2\,s(n+1)+v(n)\,s(n+2)=0, (57)
where $(u(n))_{n}$ and $(v(n))_{n}$ are two not-necessarily identical solutions of s(n+2)s(n)=0s(n+2)-s(n)=0. This includes the result of [JPNP23, Example 4.1]. Executing the following code yields the result in $0.031$ second in CPU time.
> with(NLDE:-DalgSeq); #loading the subpackage DalgSeq from NLDE
> CCfiniteToSimpleRatrec(u(n)*s(n) + 2*s(n + 1) + v(n)*s(n + 2) = 0, s(n),
[u(n + 2) - u(n) = 0, v(n + 2) - v(n) = 0], [u(n), v(n)])
$s(n+6)=\frac{s(n+5)s(n+4)s(n)-s(n+5)s(n+2)^{2}-s(n+4)^{2}s(n+1)+s(n+4)s(n+3)s(n+2)}{s(n+3)s(n)-s(n+1)s(n+2)}.$
Our CCfiniteToDalg command gives the same result (in the difference polynomial form) in $89.438$ seconds in CPU time. We provide the syntax of the code without output below to highlight the differences in syntax between the two procedures.
> CCfiniteToDalg(u*s(n) + 2*s(n + 1) + v*s(n + 2) = 0, s(n),
[u(n + 2) - u(n) = 0, v(n + 2) - v(n) = 0], [u(n), v(n)]):
Here, the $C$-finite term coefficients do not appear in the $C^{2}$-finite equation with the index variable. A full documentation of these commands is provided at https://t3gu1a.github.io/NLDEdoc/DalgSeq-Commands-and-Examples.html. $\blacksquare$
@@ 5 Subsequences
Recall that a subsequence of a sequence $(s(n))_{n}$ is a sequence $(t(n))_{n}$ such that $t(n)=s(\varphi(n))$, where $\varphi$ is a strictly increasing embedding of $\mathbb{N}$. Thus, to find a difference polynomial $p$ that vanishes at $(t(n))_{n}$, we need an endomorphism $\tilde{\sigma}$ that sends $s(\varphi(n))$ to $s(\varphi(n+1))$, i.e., the minimal gap between the orders of two difference monomials in $p$ is $\delta_{\tilde{\sigma}}=\min_{n\in\mathbb{N}}\{\varphi(n+1)-\varphi(n)\}$ and not the usual n+1n=1n+1-n=1. This section focuses on the case where $\varphi(n)=dn$. This is the $n$th term formula of an arithmetic progression of common difference $d$ with initial term 0. What makes this case rather natural is that all order gaps are multiples of the common difference $d$, and here we have $\delta_{\tilde{\sigma}}=d$. This implies that, assuming $x$ occurs in $p$, we may write $p$ as follows:
@@
The desired endomorphism is thus defined by $\tilde{\sigma}=\sigma^{d}$. The core idea of our next theorem is to exploit the flexibility in choosing the endomorphism during the construction of the dynamical system $(\mathcal{M}_{f})$ from (20). Since the sequence is assumed to be D-algebraic, no further assumptions are required to avoid zero divisors. This is because the defining difference polynomial inherently accounts for the specific indices of the subsequences, thereby ensuring the relevant denominators in the systems are nonzero.
@@
Our aim is to eliminate the difference indeterminates $y_{j}$, $j>1$. To do so, we build a radical-rational dynamical system and apply Algorithm˜1. This will yield a univariate difference polynomial in $y_{1}$ that has $(s(dn))_{n\in\mathbb{N}}$ as a zero. Let $m$ be the degree of $\sigma^{r}(x)$ in $p$ and $R_{p}$ be the rational expression obtained by solving $p=0$ for $(\sigma^{r}(x))^{m}$. Let $y_{i,j},i=1,\ldots,d$, $j=0,\ldots,r_{1}$ be new indeterminates such that
@@
The proof reduces to determining the dimension of the dynamical system obtained in the proof of Theorem˜5.1. Indeed, we have
@@
The recurrence equation of the subsequence need not be written in the same difference ring as the equation of the original sequence. In Theorem˜5.1, both $(s(n))_{n\in\mathbb{N}}$ and $(s(dn))_{n\in\mathbb{N}}$ are zeros of the constructed difference polynomial of order $dr$ in $\mathcal{D}_{\sigma}(\mathbb{K},x)$, but $(s(n))_{n}$ is generally not a zero of the corresponding difference polynomial of order $r$ in $\mathcal{D}_{\sigma^{d}}(\mathbb{K},x)$. This is illustrated in the next example.
@@
In Section˜4.1, Example˜4.7, we used the implementation from [TTW24] to convert the holonomic equation of Catalan numbers $(C(n))_{n\in\mathbb{N}}$ to the recursion
@@
One verifies that $(C(n))_{n}$ is not a solution of (76). Note that higher-order equations may be obtained here. Indeed, $(C(3n))_{n}$ is holonomic of order $1$ and degree $3$ and therefore satisfies an l.h.s. algebraic difference equation of order at most $4$ by Theorem˜4.5. Our Gröbner bases method from [TTW24] yields an equation of order $3$. $\blacksquare$
This result suggests a potential bridge between difference algebra and classical acceleration techniques. As shown in [BDGB83], the careful selection of subsequences can effectively accelerate convergence. For D-algebraic sequences, this indicates the possibility of developing acceleration techniques based entirely on algebraic computations.
@@ Acknowledgments
The author thanks Antonio Jimenez-Pastor and Veronika Pillewein for their insightful explanations of the resultant method for the D-algebraicity of $D^{n}$-finite functions, which resulted from stimulating discussions at ACA2025 (https://aca2025.github.io/) in Heraklion, Greece. The author is also grateful to Filip Mazowiecki and James Worrell for stimulating discussions regarding [CDBMP23]. This work was supported by UKRI Frontier Research Grant EP/X033813/1.